OpenStudy (anonymous):

Inequality Let f (x) =1-2x a. Find the solution set to the equation I f(x) I =5 b. Find the solution set to If(x)+1+7>/ 2

OpenStudy (callisto):

If |f(x)| = 5, then we have f(x) = 5 or f(x) = -5 ^ Solve these two equations to get the solution set for part a.

OpenStudy (callisto):

Do you mind making the second part clearer? If(x)+1+7>/ 2 seems to have something missed.

OpenStudy (anonymous):

its >with the line underneath it... get it?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

I am not sure of the left part: If(x)+1+7

OpenStudy (callisto):

You should have a pair of "|" - one is missing there.

OpenStudy (anonymous):

I f(x)+1 I >/2

OpenStudy (callisto):

Hmmm... Do you want to understand how to solve it, or do you just want to know the way to solve it?

OpenStudy (anonymous):

know the way to solve it

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

If you have |f(x)| \(\ge\) 5, then you need to solve f(x) \(\le\) -5 and f(x) \(\ge\) 5 to get the whole solution set. If you have |f(x)| \(\le\) 5, then you need to solve -5 \(\le\) f(x) \(\le\) 5 to get the solution set.